OpenStudy (anonymous):

what are the zeros of the function f(x)= x^2+x-6/x^2-x-6 a. -2,3 b. -3,2 c. -3,-2 d. 2,3

OpenStudy (anonymous):

The zeros of a rational function are the zeros of the numerator

OpenStudy (anonymous):

So long as the denominator isn't also zero

OpenStudy (anonymous):

@wio so it would be -6?

OpenStudy (anonymous):

no, because it doesn't equal zero at that point

OpenStudy (anonymous):

none of the answer choices quals zero.

OpenStudy (anonymous):

\[ (2)^2+(2)-6 = 4-4 = 0 \]

OpenStudy (anonymous):

So \(2\) is a zero

OpenStudy (anonymous):

oh i get it now.

OpenStudy (anonymous):

plug in the answer choice in the equation :)

OpenStudy (anonymous):

and see which equals 0

OpenStudy (anonymous):

thanks for the help!

OpenStudy (radar):